When 60.8 g of calcium is reacted with nitrogen gas, 28.7 g of calcium nitride is produced.

What is the percent yield of calcium nitride for this reaction?
3Ca(s)+N2(g)→Ca3N2(s)
Express your answer with the appropriate units.
1
Expert's answer
2015-10-19T10:53:21-0400
Percent yield = actual yield / theoretical yield * 100%
Theoretical yield (Ca3N2) = n (Ca3N2) * M (Ca3N2) = m (Ca) / M (Ca) * 3 * M (Ca3N2) = 60, 8 g / 40 g/mol / 3 * 148 g / mol = 75 g
Percent yield (Ca3N2) = 28,7 g / 75 g * 100% = 38,3 %.